# Soft-Clamped Perimeter Modes of Polygon Resonators

### v1

@@ Abstract
To investigate the hybridization between the bending modes of the polygon sides and the torsional modes of the tethers, we set $l_{\text{s}}=700~{}\upmu\mathrm{m}$ and varied the length of the tethers while keeping all other parameters constant. Fig. 4 plots the dissipation dilution factor as a function of the tether length in the polygon resonator, revealing three well-defined resonant peaks. The initial peak exhibits the maximum dissipation dilution, while subsequent peaks display a progressive reduction in magnitude. This behavior originates from the torsional angle at the tether-polygon junction modulating the transverse displacement derivative of the polygon side - a quantity whose squared value directly governs the tensile energy based on Eq. 8. Crucially, the observed first peak position occurs at $kl_{t}\approx\pi/2$, whereas the trough coincides with $kl_{t}\approx\pi$, consistent with the theoretical model according to Eq. 14. These results confirm that tailored adjustment of tether length permits precise control over the quality factors of the resonator and offers valuable guidance to optimize high-performance polygon resonators. As the tether length increases, we also noticed that the peak positions predicted by the analytical model exhibit a growing deviation from the FEM simulations. This physical mechanism underlying the increase in $k$ with $l_{\text{t}}$ remains under investigation.
@@
An important issue in the study of polygon structures involves the enhancement of dissipative dilution in the supporting tethers. This phenomenon has been demonstrated in doubly-clamped stressed nanoribbons, similar to that of supporting tethers in polygon structures (see the supplementary material for a detailed derivation). Thus, if only the tethers are considered in a polygon structure system, the dissipation factor for the $n^{\text{th}}$ torsional mode of the tethers can be expressed as follows:
@@
The dimensions of the tethers ($l_{t}\gg w_{t}$) and material properties of the silicon nitride film (see the simulation part) suggest that the first term $\frac{2G}{\sigma_{t}}\frac{h^{2}}{w_{t}^{2}}$ in the denominator in Eq. 15 governs energy dissipation. This finding aligns with prior work Pratt et al. (2023) and reflects a limited torsional dissipation dilution.
@@
In this work, we present a comprehensive analytical study of perimeter vibrational modes in polygonal resonators. We derive the Timoshenko-Gere equation under the high tension condition for the angular displacement of nanoribbons. The modified equation is extended to the torsional vibration of supporting thethers of polygon resonators and accurately predict both the resonant frequencies and dissipation dilution factors of perimeter modes. The quality factor of the fundamental mode enhances by two orders of magnitude compared to uniform strings of the same dimensions. Moreover, our analysis reveals that the resonance of the tethers can significantly degrade the dissipation dilution factor of the perimeter mode, and the tether length can be used to tune the dilution factor. Our findings provide critical insights for designing high-$Q$ polygon resonators, with direct applications in cavity optomechanics Aspelmeyer et al. (2014). Additionally, our work impacts nanoscale force sensing Moser et al. (2013); Gavartin et al. (2012), enabling precision measurements in weak-force detection. Furthermore, our work advances the study of multimode phononic structures Faust et al. (2012); Peano et al. (2015); Lu et al. (2017), offering new avenues for exploring topological phononics and non-Hermitian wave control.

### v2

@@ Abstract
To investigate the hybridization between the bending modes of the polygon sides and the torsional modes of the tethers, we fix $l_{\text{s}}=700~\upmu\mathrm{m}$ and vary the length of the tethers while keeping all other parameters constant. Fig. 4 plots the dissipation dilution factor as a function of the tether length in the polygon resonator, revealing three well-defined resonant peaks. The first peak yields the maximum dissipation dilution, while subsequent peaks exhibit progressively reduced magnitudes. This behavior originates from the torsional rotation at the tether-polygon junction, which modulates the slope of the transverse displacement of the polygon side - a quantity whose squared value directly governs the tensile energy based on Eq. 9. Consistent with the model, both the first peak at $kl_{t}\approx\pi/2$ and the first trough at $kl_{t}\approx\pi$ are in excellent agreement with Eq. 15. The torsion-angle profiles at $kl_{t}\approx\pi$, which correspond to tether lengths of 3920, 7890, and $11810~\upmu\mathrm{m}$, are further visualized in Fig. 4. These results indicate that adjustment of the tether length allows precise control over the quality factors of the resonator and offers valuable guidance to optimize high-performance polygon resonators. As the tether length increases, we also notice that the peak positions predicted by the analytical model exhibit a growing deviation from the FEM simulations. This physical mechanism underlying the increase in $k$ with $l_{\text{t}}$ remains under investigation.
While our theoretical model accounts for distributed bending loss in the polygon sides and torsional loss in the tethers, it neglects the torsional loss in the polygon sides; this simplification could also cause the discrepancy between the analytical model and FEM simulation, especially for wider designs. Analogous to neglecting the boundary loss of sufficiently long polygon sides, we can also define a range for the polygon side width $w_{s}$ within which its torsional loss is negligible. To validate this, we perform simulations with $l_{s}=l_{t}=700~\upmu\mathrm{m}$ and $w_{t}=400~\mathrm{nm}$, and scan $w_{s}$ from $100~\mathrm{nm}$ to $2000~\mathrm{nm}$. As shown in Fig. 4, the dissipation dilution factor increases monotonically as the width increases from $100~\mathrm{nm}$ to $400~\mathrm{nm}$, suggesting that the torsional loss is negligible in this regime. For widths from $400~\mathrm{nm}$ to $2000~\mathrm{nm}$, the close agreement between our theoretical predictions and the FEM results, with a deviation of merely 6%, further confirms that the polygon side torsional loss has an insignificant impact across the entire studied range.
To systematically explore the impact of resonator dimensions on the quality factor and identify the optimal configuration, we employ a parametric approach, by independently varying tether and polygon side dimensions. First, we fix $l_{s}=700~\upmu\mathrm{m}$, $w_{s}=300~\mathrm{nm}$, while scanning the tether length $l_{t}$ over a range of $1-4000~\upmu\mathrm{m}$ and the tether width $w_{t}$ over $1-2000~\mathrm{nm}$. The results, as shown in Fig. 4, are consistent with those in Fig. 4. Several local maxima of $D_{Q}$ are observed, with the highest value occurring at point ($1700~\upmu\mathrm{m}$, $450~\mathrm{nm}$). In Fig. 4, with the tether length and width fixed at $700~\upmu\mathrm{m}$ and $400~\mathrm{nm}$, respectively, $D_{Q}$ increases monotonically with $l_{s}$. This is primarily driven by the $D_{Q}\propto l_{s}^{2}$ dependence originating from the $\lambda_{l}^{2}$ term in Eq. 14 and Eq. 15. Moreover, the optimal $w_{s}$ increases with $l_{s}$, which can be attributed to the linear dependence of the $kl_{s}/[(1+\nu)r_{w}]$ term in Eq. 15 on $l_{s}$.
Our analytical model provides a reliable estimate of the dissipation dilution factor $D_{Q}$ for perimeter modes, accurately capturing its dependence on the mode number, polygon side width, and tether length. As shown in Fig. 4, the model agrees well with FEM simulations for short and moderate long tethers. The noticeable discrepancy emerges in the limit of extremely long tethers or excessively wide side widths, which can be attributed to a simplified assumption in the model, namely that the axial stress in the tethers remains constant while the tether length varies. Prior studies Bückle et al. (2021) have indicated that increasing the tether length reduces the axial stress. Within our model, this stress relaxation modifies the effective torsional stiffness, as described in Eq. 7, leading to a weak dependence of $k$ on $l_{t}$ (Eq. 8) and a corresponding shift in the resonance peaks. Meanwhile, numerical convergence challenges in FEM for extremely high-aspect-ratio geometries cannot be excluded, although they are not expected to dominate the noticeable discrepancy. It is worth noting that, while the effect of the torsional loss in the polygon sides is neglected here, the side width corresponding to the maximum $D_{Q}$ for a fixed polygon side length lies within a range where the torsional effect in the polygon sides is negligible.
@@
An important issue in the study of polygon structures involves the enhancement of dissipation dilution in the supporting tethers. This phenomenon has been demonstrated in doubly-clamped stressed nanoribbons, which are analogous to the supporting tethers in polygon structures (see Appendix B for a detailed derivation). Thus, considering the tethers within a polygon resonator system, the dissipation dilution factor for the $n^{\text{th}}$ torsional mode is given by:
@@
The dimensions of the tethers ($l_{t}\gg w_{t}$) and material properties of the Si3N4 film (see the simulation part) suggest that the first term $\frac{2G}{\sigma_{t}}\frac{h^{2}}{w_{t}^{2}}$ in the denominator of Eq. 16 dominates the energy dissipation. This finding aligns with prior work Pratt et al. (2023) and points to a limitation in torsional dissipation dilution.
@@
In this work, we present a comprehensive analytical study of perimeter vibrational modes in polygonal resonators. We derive the TG equation for the angular displacement of nanoribbons under the high tension condition. The modified framework can be extended to the torsional vibration of supporting tethers of polygon resonators, accurately predicting both the resonance frequencies and dissipation dilution factors of perimeter modes. Notably, the quality factor of the fundamental mode is enhanced by two orders of magnitude compared to uniform strings of the same dimensions as the polygon sides. Our analysis reveals that the resonance of the tethers can significantly degrade the dissipation dilution factor of the perimeter mode, and the tether length can be used to tune the dilution factor. Our findings reveal that the dissipation dilution factor of the perimeter mode in polygon resonators scales as $1/\lambda^{2}$, consistent with the existing literature Bereyhi et al. (2022a). This scaling law provides a transformative strategy for designing ultra-high-$Q$ resonators, with direct applications in cavity optomechanics Aspelmeyer et al. (2014). Concurrently, we elucidate the torsional dissipation dilution limit of the supporting tethers. Additionally, our work has significant implications for nanoscale force sensing Moser et al. (2013); Gavartin et al. (2012), enabling precision measurements in weak-force detection. Furthermore, our work advances the study of multimode phononic structures Faust et al. (2012); Peano et al. (2015); Lu et al. (2017), offering new avenues for exploring topological phononics and non-Hermitian wave control.
@@ I Acknowledgements
Appendix A Appendix A: Free Torsional Vibration Under Tensile Stress
The theoretical foundation of torsional dissipation dilution can be traced back to Buckley’s pioneering work on twisted elastic strips Buckley (1914), which demonstrates that tensile stress can geometrically enhance torsional stiffness without proportionally increasing dissipative losses. This phenomenon arises because tensile stress preferentially aligns the principal strain axes along the ribbon’s length, thereby suppressing transverse deformation through geometric stiffening, while intrinsic loss mechanisms remain primarily governed by material properties Trahair (2017) (p258-265). Buckley’s analysis of bifilar systems reveals that torsion modes of ribbons are naturally soft-clamped Tsaturyan (2017) and the effective quality factor $Q_{\mathrm{eff}}$ can be enhanced beyond the material limit through stress engineering, a critical insight for modern nanomechanical resonators Fedorov et al. (2019).
The study by Pratt et al. Pratt et al. (2023) significantly extends Buckley’s paradigm by demonstrating experimentally the validation of extreme dissipation dilution ($Q\cdot f>10^{13}$ Hz) in nanoribbons made of $\text{Si}_{3}\text{N}_{4}$ at room temperatures. Their work establishes a novel theoretical framework for quantum precision measurements exploiting the torsion modes of nanomechanical resonators Kim et al. (2016); Enomoto et al. (2016); Treps et al. (2003), enabling novel designs where torsional modes benefit from dissipation dilution Bereyhi et al. (2022a); Ghadimi (2018).
Here, we derive the general solution for the torsional vibrational modes of a stressed nanoribbon based on Hamilton’s variational principle. First, we revisit the TG theory for torsional vibration of ribbons in the cartesian representation Rao (2019) (p296-302). For a normal nanoribbon resonator with length $L$, width $w$, and thickness $h$ satisfying $L>w>h$, the equation of motion for the angular displacement $\Theta(x,t)$ is
$\begin{split}\rho I_{p}\frac{\partial^{2}\Theta}{\partial t^{2}}-\frac{\partial^{2}}{\partial x\partial t}(\rho I_{\psi}\frac{\partial^{2}\Theta}{\partial x\partial t})&-\frac{\partial}{\partial x}(C\frac{\partial\Theta}{\partial x})\\ &+\frac{\partial^{2}}{\partial x^{2}}(EI_{\psi}\frac{\partial^{2}\Theta}{\partial x^{2}})=0,\end{split}$ (A1)
where $x$ denotes the axial coordinate along the ribbon, $I_{p}=\iint_{A}(y^{2}+z^{2})\mathrm{d}A$ is the geometric moment of inertia, $I_{\psi}=\iint_{A}\psi^{2}\mathrm{d}A$ ($\psi$ $\approx-yz$ is the warping function Chopin and Filho (2019)), $A=w\cdot h$ the cross section, $C$ the torsional rigidity ($C=\iint_{A}G[(\frac{\partial\psi}{\partial y}-z)^{2}+(\frac{\partial\psi}{\partial z}+y)^{2}]\mathrm{d}A$), $E$ the Young’s modulus, $G$ the shear modulus, and $\rho$ the material density. In the case of a highly stressed ribbon, the angular displacement is obtained by minimizing the action integral $\mathcal{S}=\int_{t_{1}}^{t_{2}}(T-W_{\text{strain}}-W_{\text{tens}})\,dt$, where $T$, $W_{\text{strain}}$ and $W_{\text{tens}}$ represent the kinetic, elastic strain and tensile energies, respectively. The modified TG equation can be written as
$\begin{split}\rho I_{p}\frac{\partial^{2}\Theta}{\partial t^{2}}-\frac{\partial^{2}}{\partial x\partial t}(\rho I_{\psi}\frac{\partial^{2}\Theta}{\partial x\partial t})&-\frac{\partial}{\partial x}((C+\sigma I_{y})\frac{\partial\Theta}{\partial x})\\ &+\frac{\partial^{2}}{\partial x^{2}}(EI_{\psi}\frac{\partial^{2}\Theta}{\partial x^{2}})=0,\end{split}$ (A2)
where $I_{y}$ is the geometric moment of inertia with respect to the $y$-axis, and $\sigma=\sigma_{xx}$ is the axial stress component. By applying separation of variables, the solution to this differential equation takes the form of a superposition of normal modes, each consisting of a position-dependent component multiplied by a time-dependent component:
$\Theta(x,t)=\sum_{n=1}^{\infty}u_{0,n}\cos{(\omega_{n}t)}\theta_{n},$ (A3)
where $\omega_{n}$ is the eigenfrequency of the $n^{\text{th}}$ mode, $\theta_{n}$ the mode shape, and $u_{0,n}$ the amplitude. Inserting Eq. A3 into Eq. A2 yields
$EI_{\psi}\frac{\partial^{4}\theta_{n}}{\partial x^{4}}+(\rho I_{\psi}\omega_{n}^{2}-C-\sigma I_{y})\frac{\partial^{2}\theta_{n}}{\partial x^{2}}-\rho I_{p}\omega_{n}^{2}\theta_{n}=0.$ (A4)
Assuming a sinusoidal mode shape with wavenumber $\beta_{n}$, and applying the boundary conditions $\theta_{n}(0)=\theta_{n}(L)=\frac{\partial^{2}\theta_{n}}{\partial x^{2}}|_{0}=\frac{\partial^{2}\theta_{n}}{\partial x^{2}}|_{L}=0$ yield
$\theta_{n}(x)=\sin{(\beta_{n}x)},~\beta_{n}=\frac{n\pi}{L}.$ (A5)
Combining Eq. A5 with Eq. A4 gives
$\begin{split}\rho w^{2}h^{2}\beta_{n}^{2}\omega_{n}^{2}+12\rho(w^{2}+h^{2})\omega_{n}^{2}&=Ew^{2}h^{2}\beta_{n}^{4}\\ &+48Gh^{2}\beta_{n}^{2}+12w^{2}\sigma\beta_{n}^{2}.\end{split}$ (A6)
The first term $\rho w^{2}h^{2}\beta_{n}^{2}\omega_{n}^{2}$ is negligible due to $\rho w^{2}h^{2}\beta_{n}^{2}\ll 12\rho(w^{2}+h^{2})$. Similarly, the term $Ew^{2}h^{2}\beta_{n}^{4}$ exhibits negligible influence on the system dynamics compared to the two dominant terms dominating the right-hand side. The mode frequency $\omega_{n}$ can be written in the following form
$\begin{split}\omega_{n}&=\sqrt{\frac{48Gh^{2}\beta_{n}^{2}+12w^{2}\sigma\beta_{n}^{2}}{12\rho(w^{2}+h^{2})}}\\ &\approx\frac{n\pi}{L}\sqrt{\frac{\sigma}{\rho}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}.\end{split}$ (A7)
The full set of wave numbers can be found by solving the dispersion relation of the pre-stressed ribbon:
$\beta_{n}^{4}+(\frac{48G}{Ew^{2}}+\frac{12\sigma}{Eh^{2}})\beta_{n}^{2}-\frac{12\rho}{Eh^{2}}\omega_{n}^{2}=0.$ (A8)
Four solutions of $\beta_{n}$ can be found:
$\beta_{n,1-2}=\beta_{\sigma}=\frac{n\pi}{L},$ (A9)
$\beta_{n,3-4}=\beta_{E}=\sqrt{\frac{12\sigma}{Eh^{2}}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}.$ (A10)
Within the analytical model, $\beta_{\sigma}$ is the already known wavenumber of a perfect ribbon, and $\beta_{E}$ is the wave number related to the torsional stiffness of the ribbon. Under clamped boundary conditions, the mode shape takes the following form:
$\theta_{n}(x)=\sin{(\beta_{\sigma}x)}+\frac{\beta_{\sigma}}{\beta_{E}}[e^{-\beta_{E}x}-\cos{(\beta_{\sigma}x)}].$ (A11)
This equation enables the decoupling of the idealized torsion mode from the boundary-constrained components in the actual vibrational mode. It reveals that $\beta_{E}$ corresponds to the inverse characteristic length of the exponentially decaying edge effect.
Appendix B Appendix B: Torsional Dissipation Dilution of Nanoribbons
The quality factor of a stressed nanomechanical resonator exhibits significant enhancement due to dissipation dilution. For a ribbon resonator under tensile stress $\sigma$, the dissipation dilution factor $D_{Q}$ can be written as
$D_{Q,n}=\frac{Q}{Q_{\text{int}}}=1+\frac{\langle W_{\text{tens}}\rangle}{\langle W_{\text{lossy}}\rangle}$}, (B1)
where $\langle W_{\text{tens}}\rangle$ is the energy stored in the stress potential, and $\langle W_{\text{lossy}}\rangle$ is the energy concentrated in the bending and shear components Fedorov et al. (2019); Sementilli et al. (2022). The dissipation dilution factor $D_{Q}$ measures the proportional enhancement in the mechanical quality factor resulting from tensile stress. To analytically estimate $D_{Q}$, we express a lossless tensile energy based on Eq. A3:
$\langle W_{\text{tens}}\rangle=\frac{\sigma}{2}\frac{hw^{3}}{12}\int u_{0,n}^{2}\langle(\frac{\partial\theta_{n}}{\partial x})^{2}\rangle dx.$ (B2)
The lossy energy is
$\langle W_{\text{lossy}}\rangle=\frac{E}{2}\frac{wh^{3}}{12}\int u_{0,n}^{2}\langle\frac{2}{1+\nu}(\frac{\partial\theta_{n}}{\partial x})^{2}+\frac{w^{2}}{12}(\frac{\partial^{2}\theta_{n}}{\partial x^{2}})^{2}\rangle dx.$ (B3)
Substituting Eq. B2 and Eq. B3 into Eq. B1 yields a dissipation dilution factor of
$D_{\text{Q},n}=1+\frac{1}{\frac{2G}{\sigma}\frac{h^{2}}{w^{2}}+\frac{n^{2}\pi^{2}}{12}\frac{Eh^{2}}{\sigma L^{2}}+\frac{h}{L}\sqrt{\frac{E}{3\sigma}(1+\frac{4G}{\sigma}\frac{h^{2}}{w^{2}})}}.$ (B4)
The first two terms in the denominator arise from distributed shear deformation and distributed bending, respectively, while the last term is due to the coupled bending-shear interaction at the clamps. For a standard long-string resonator, the length is much greater than the thickness, causing the first term to dominate Sadeghi et al. (2020). In commercial high-stress silicon nitride, the ratio of Young’s modulus to stress is approximately 250 Fedorov et al. (2020), leading to a calculated $D_{Q}$ of $\sim 1$. Consequently, the torsion mode cannot experience dissipation dilution Fedorov et al. (2019). In contrast, for a nanoribbon, the width substantially exceeds the thickness as well, resulting in Q factors which scale as $(w/h)^{2}$ as high as $10^{8}$ Pratt et al. (2023). Provided that curvature at the clamping points is nonnegligible, Eq. B4 can be reduced to
$D_{Q,n}\approx 1+\frac{1}{\frac{2G}{\sigma}\frac{h^{2}}{w^{2}}+\frac{n^{2}\pi^{2}}{12}\frac{Eh^{2}}{\sigma L^{2}}+\frac{h}{L}\sqrt{\frac{E}{3\sigma}}}.$ (B5)
The formulation bears strong similarity to the dissipation dilution factor expression for standard string resonators Schmid et al. (2016); Sadeghi et al. (2019). This correspondence suggests that established techniques for enhancing the quality factor in flexural modes of string resonators - including but not limited to phononic crystal engineering Beccari et al. (2022), tapered clamping structures Bereyhi et al. (2019), and hierarchical mechanical designs Bereyhi et al. (2022b); Fedorov et al. (2020)- can be effectively extended to enhance the torsion mode performance in ribbon resonators.